OpenStudy (anonymous):

SOMEONE HELP ME PLSSS :D which rational number lies between -7/8 and 0.3?? -1/8, 1/2, 0.75, -0.9

OpenStudy (lgbasallote):

-7/8 = -0.875 so find a number between -0.875 and 0.3 -1/8 = -0.125 1/2 = 0.5 0.75 = 0.75 -0.9 = -0.9 so what do you think is -0.875<x<0.3

OpenStudy (experimentx):

1/8

OpenStudy (experimentx):

oops sorry -1/8

OpenStudy (anonymous):

thanks